OpenStudy (anonymous):

Let C be the positively oriented square with vertices (0,0) , (12,0) , (12,1) , (0,1) . Use Green's Theorem to evaluate the line integral CFdr where F(xy)=8e^yi+7xe^yj

OpenStudy (anonymous):

nm got it

OpenStudy (anonymous):

i think its weird to get a negative answer if the parameters are based in the positive first quadrant

OpenStudy (anonymous):

can u post ur solution?

OpenStudy (anonymous):

F(x,y)= 8e^yi+7xe^yj the line integral \[\int\limits_{?}^{?}F.dr\]

OpenStudy (anonymous):

F(x,y)=<8e^y,7e^y>, where P= 8e^y and Q= 7e^7, Qx=7e^y and Py=8e^y. Greens theorem states int(int(Qx-Py,,dx)dy)

OpenStudy (anonymous):

0<x<12 and 0<y<1

OpenStudy (anonymous):

=\[\int\limits_{?}^{?}F(r(t)),r'(t) dr\]

OpenStudy (anonymous):

thats the fundamental theorem for conservative line intergrals

OpenStudy (anonymous):

i m trying to use the line integral for a vector field

OpenStudy (anonymous):

int(int(7e^y-8e^y,y,0,1),x,0,12)=-20.619

OpenStudy (anonymous):

:S

OpenStudy (anonymous):

y negative?

OpenStudy (anonymous):

I guess the function is loopy

OpenStudy (anonymous):

haha